### v2

# EuleroDec: A Complex-Valued RVQ-VAE for Efficient and Robust Audio Coding

Abstract
Audio codecs power discrete music generative modelling, music streaming and immersive media by shrinking PCM audio to bandwidth-friendly bit-rates. Recent works have gravitated towards processing in the spectral domain; however, spectrogram-domains typically struggle with phase modeling which is naturally complex-valued. Most frequency-domain neural codecs, either disregard phase information or encode it as two separate real-valued channels, limiting spatial fidelity. This entails the need to introduce adversarial discriminators at the expense of convergence speed and training stability to compensate for the inadequate representation power of the audio signal. In this work we introduce an end-to-end complex-valued RVQ‑VAE audio codec that preserves magnitude–phase coupling across the entire analysis–quantization–synthesis pipeline and removes adversarial discriminators and diffusion post‑filters. Without GANs or diffusion we match or surpass much longer-trained baselines in-domain and reach SOTA out-of-domain performance. Compared to standard baselines that train for hundreds of thousands of steps, our model reducing training budget by an order of magnitude is markedly more compute-efficient while preserving high perceptual quality.
Index Terms—  Neural Audio Codecs, EuleroDec, Vector-quantized variational autoencoders (VQ-VAE), Complex-Valued Neural Networks
1 Introduction
Spectral-domain audio codecs decompose a signal into its time-frequency representation, typically via short-time Fourier transform (STFT). Although the magnitude (or power) spectrum carries most perceptual content, the phase spectrum is equally essential for faithful reconstruction: improper or discarded phase information leads to the smearing of transients, ’phasiness’, and audible artifacts in the decoded signal. Many recent approaches, both in the waveform and spectral domains, employ multi-scale adversarial discriminators [defossez2022high] [kumar2023high] or score or flow-based post filters [welker2025flowdecflowbasedfullbandgeneral] [wu2024scoredec] [10890277] to improve perceptual quality at the price of adversarial instability and slower convergence. An alternative line of work encodes complex spectra by splitting real-imaginary or magnitude–phase components and processing them independently [wu2025complexdec, du2024apcodecspectrumcodingbasedhighfidelityhighcompressionrate, ai2024apcodecneuralaudiocodec, du2023funcodecfundamentalreproducibleintegrable]. However, this factorization neglects intrinsic phase–magnitude coupling and leads to parameter redundancy. Complex-valued neural networks (CVNNs) have shown promise in speech enhancement [hu2020dccrndeepcomplexconvolution, halimeh2021complexvaluedspatialautoencodersmultichannel] and generative modeling [nakashika20_interspeech]. Yet existing codec-oriented approaches either operate partially in the real domain [Ru_2024], breaking phase coherence, or are restricted to music and a single bitrate. To date, to the best of our knowledge, no codec realizes a fully end-to-end complex-valued analysis-quantization-synthesis pipeline without adversarial or post-filtering components. In this work, we introduce the first fully end-to-end complex-valued neural audio codec for 24 kHz speech at 6 and 12 kbps. All stages of the pipeline - from waveform input to waveform reconstruction - are complex-valued and jointly optimized using Wirtinger calculus, without real-valued detours. The model integrates complex convolutions, attention, and normalization with 2×2 whitening, preserving the algebraic structure of the STFT and intrinsic amplitude–phase coupling. Without adversarial training or diffusion-based post-filters, the proposed approach achieves consistent in-domain and out-of-domain improvements at the target bitrates. Our contributions are as follows: (1) We propose the first end-to-end complex-valued VQ-VAE for audio, (2) our codec attains state-of-the-art quality without adversarial discriminators or score-based post-filters; (3) we reach and surpass state-of-the-art quality with fast, steady convergence and reduces training budget by $95\%$.
2 Background and Related Work
In this section, we give a brief introduction to Residual Vector Quantization [lee2022autoregressiveimagegenerationusing] and complex-valued neural networks.
2.1 Residual Vector Quantization
In an audio codec based on an RVQ–VAE [ai2024apcodecneuralaudiocodec] [du2024apcodecspectrumcodingbasedhighfidelityhighcompressionrate] [wu2025complexdec], the encoder projects the complex STFT frame (or the waveform when working directly in the time-domain [defossez2022high] [borsos2023soundstormefficientparallelaudio][du2023funcodecfundamentalreproducibleintegrable] [kumar2023high]) into a latent vector $\mathbf{z}\!\in\!\mathbb{R}^{H}$. A stack of $M$ residual codebooks $\{\mathcal{E}^{(m)}\}_{m=1}^{M}$ iteratively approximates $\mathbf{z}$:
$\mathbf{r}^{(m)}=\mathbf{z}-\!\!\sum_{j<m}\mathbf{e}_{k_{j}}^{(j)},\qquad k_{m}=\operatorname*{arg\,min}_{k}\bigl\lVert\mathbf{r}^{(m)}-\mathbf{e}_{k}^{(m)}\bigr\rVert_{2},$
where $\mathbf{e}_{k}^{(m)}$ is the selected centroid at stage $m$. Only the index sequence $(k_{1},\dots,k_{M})$ is transmitted, yielding a bit-rate $R_{f}\,M\,\log_{2}K$. Because the latent space is real, existing spectral codecs must split the complex STFT into two real signals before quantisation. The first prevalent strategy [ai2024apcodecneuralaudiocodec] [du2024apcodecspectrumcodingbasedhighfidelityhighcompressionrate] [feng2025stftcodechighfidelityaudiocompression] writes the spectrum as modulus $|X|$ and unwrapped phase $\angle X$; two independent RVQ pipelines are then trained, typically with a log-power loss for magnitude and a wrapped angular loss for phase.
Fig. 1: Model Architecture (All layers are complex-valued)
The second strategy [wu2025complexdec] [feng2025stftcodechighfidelityaudiocompression] separates the spectrum into its real and imaginary components $\Re\{X\}$ and $\Im\{X\}$, feeds each map to an RVQ cascade optimised by Euclidean mean-squared error, and recombines the output as $\hat{X}=\hat{R}+j\hat{I}$ prior to the inverse STFT. Both approaches, while duplicating the entire RVQ path into different streams, neglect the intrinsic correlation between magnitude and phase.
2.2 Complex-valued neural networks
Complex-valued neural networks [DBLP:journals/corr/TrabelsiBSSSMRB17] represent inputs, weights, and activations as $z=x+iy$ and compute with true complex algebra. A complex convolution is linear over $\mathbb{C}$ and couples real and imaginary parts:
$(w*z)[n]$ $=\sum_{k}w_{k}\,z_{n-k},$
$w_{k}$ $=a_{k}+ib_{k},\qquad z_{n-k}=x_{n-k}+iy_{n-k}.$
This coupling lets the model learn amplitude-phase interactions rather than treating $x$ and $y$ as independent channels. A key property is phase equivariance [DBLP:journals/corr/TrabelsiBSSSMRB17]: for any $\phi\in\mathbb{R}$,
$f\!\left(e^{i\phi}z\right)\;=\;e^{i\phi}f(z),$
which preserves the geometry induced by $U(1)$ rotations. Nonlinear layers are chosen to respect this symmetry. The modReLU applies a threshold to the modulus, leaving the phase intact,
$\mathrm{modReLU}(z)\;=\;\mathrm{ReLU}(|z|+b)\,\frac{z}{|z|}.$
Fig. 2: Visualization of the modReLU activation: it applies a threshold to the modulus, leaving the phase intact.
The effect of this function is in Figure. 2 Normalization models cross-channel dependence by whitening $(x,y)$ jointly with a $2\times 2$ covariance, rather than normalizing parts separately. In audio coding, where phase carries timing and interference patterns, these properties make CVNNs a natural fit for spectrogram processing and improve generalization beyond the training domain [Eilers_2023].
3 METHODS
We work directly in the complex domain with
$x\in\mathbb{C}^{B\times C\times F\times T}$, of type complex64. We compute complex spectrograms with STFT at $24\,\text{kHz}$ $(N_{\text{FFT}}=512,\ \text{win}=512,\ \text{hop}=64,\ \text{Hann})$ over 256 frames. For coding, we use RVQ with $2048$-entry codebooks. At $6\,\text{kbps}$, a temporal stride of $8$ reduces 256 frames to 32 latent frames ($\approx 46.9\ \text{tokens/s}$). With fixed-length coding (11bits/symbol11\ $\text${bits/symbol} for $K=2048$) and $12$ codebooks, this yields a bitrate of $\approx 6.2\,\text{kbps}$. At $12\,\text{kbps}$, we use a temporal stride of $4$, doubling the token rate to $\approx 93.8\ \text{tokens/s}$ while keeping the same number of codebooks, resulting in a bitrate of $\approx 12.4\,\text{kbps}$. As shown in  1, the architecture is a complex-valued VQ-VAE. All layers are complex-native: complex convolutions, complex normalizations (complex BatchNorm and complex RMSNorm), complex activations (ModReLU or GELU), and complex axial attention [Eilers_2023].
3.1 Encoder and Decoder
[…]
Stage $C_{\text{out}}$ Kernel Stride Padding
1 48 (6,6) (2,2) (2,2)
2 64 (6,1) (2,1) (2,0)
3 96 (4,4) (2,2) (1,1)
4 128 (4,4) (2,2) (1,1)
The encoder begins with five complex residual layers using dilations [yu2017dilatedresidualnetworks] ((1,1),(3,3),(3,5),(3,7),(1,1))((1,1),\,(3,3),\,(3,5),\,(3,7),\,(1,1)) to enlarge the receptive field while maintaining stable complex statistics. A complex $3\times 7$ convolution then prepares features for hierarchical compression. Four downsampling stages follow. In each stage, a gated skip branch computes adaptive complex average pooling [DBLP:journals/corr/TrabelsiBSSSMRB17] of the input and applies a $1\times 1$ complex projection; this branch is summed with a main path that performs complex downsampling, complex normalization [abdalla2023complexvaluedneuralnetworks], complex activation, a $3\times 3$ complex convolution, complex axial self-attention [Eilers_2023] along time, another $3\times 3$ complex convolution, and a final $1\times 1$ complex projection. The strided branch is summed with drop-path probability $p=0.05$.
Fig. 3: UMAP 2D of encoder embeddings and stage-0 centroids at 6 kbps. Our RVQ shows 100 % code utilization and an effective perplexity of 73.2% of the available codes, indicating broad, non-collapsed usage (a collapsed codebook would yield low perplexity). Distances are qualitative
We keep the $2\mathrm{D}$ spectrogram structure throughout the encoder so that frequency bins retain spatial relations rather than being collapsed into channels at early stages. Once spatial scales are compressed, the model applies axial attention across frequency and a complex feed-forward block before the quantizer. The decoder mirrors this mechanism without using the pooling branch, starting with frequency-axis attention and a complex feed-forward block, then four upsampling stages that invert the encoder ordering and restore the full-resolution complex spectrogram.
3.2 Vector Quantizer
We remain in the complex domain throughout quantization. The encoder output $z_{e}\in\mathbb{C}^{B\times C\times F\times T}$ is reshaped by collapsing frequency into channels, yielding $z_{e}^{\flat}\in\mathbb{C}^{B\times(C\!\cdot\!F)\times T}$. A complex linear projection $W_{\text{in}}\in\mathbb{C}^{D\times(C\cdot F)}$ maps this merged representation to the code dimension. We then apply a residual vector quantizer with $S$ stages $(S=12\text{ at }6\,\mathrm{kbps})$. Codebooks are initialized after a brief warm-up of 30 optimization steps by sampling centroid seeds from the current continuous encoder embeddings and adding small complex Gaussian noise, which increases diversity at start-up while keeping centroids near the data manifold. At each stage, for every time index, vector quantization selects the nearest complex centroid under the Hermitian-induced Euclidean metric. Given $\mathcal{E}=\{e_{k}\}_{k=1}^{K}\subset\mathbb{C}^{D}$ and $x\in\mathbb{C}^{D}$,
$d_{k}(x)$ $=\lVert x\rVert_{2}^{2}+\lVert e_{k}\rVert_{2}^{2}-2\,\mathrm{Re}\!\left(x^{\mathrm{H}}e_{k}\right),$
$k^{\star}(x)$ $=\arg\min_{k}d_{k}(x).$
The stage output accumulates a quantized reconstruction and updates the residual passed to the next stage. Encoder stability is promoted by a commitment loss that pulls $z_{e}$ toward its assigned centroid without moving the codebook:
$\mathcal{L}_{\text{commit}}\;=\;\beta\,\frac{1}{N}\sum_{n=1}^{N}\left\|\,z_{e,n}-\operatorname{sg}\!\big(e_{k^{\star}(n)}\big)\right\|_{2}^{2}.$
Where $sg$ is the stop gradient operation. Codebooks are updated via exponential moving averages of assignment counts and feature sums. To prevent early stickiness, we warm up the EMA [Lancucki_2020] decay $m$ across epochs, starting smaller ($0.98$) and annealing to the target value ($0.999$), which increases the effective codebook learning rate at the beginning and stabilizes it later.